OpenStudy (anonymous):

show that y = (2/3)e^x + e^(-2x) is a solution of the differential equation y' +2y = 2e^x

OpenStudy (anonymous):

I'm stuck because I can't get x and dx; and y and dy together by division and multiplication.

myininaya (myininaya):

Do you know how to differentiate y w.r.t. x ?

myininaya (myininaya):

\[\frac{d}{dx}(y)=\frac{d}{dx}(\frac{2}{3}e^{x}+e^{-2x}) \\ \frac{dy}{dx}=\frac{d}{dx}\frac{2}{3}e^x+\frac{d}{dx}e^{-2x}\]

myininaya (myininaya):

can you find the derivative of (2e^x/3) and e^(-2x)?

OpenStudy (anonymous):

yes

myininaya (myininaya):

ok then when you are done finding y' there ... plug it into y'+2y

myininaya (myininaya):

and you want to see if y'+2y will simplify to 2e^x

myininaya (myininaya):

@yoonlim19 are you still having problems on this question?

OpenStudy (anonymous):

I GOT IT :D Thak you :)

myininaya (myininaya):

np

OpenStudy (solomonzelman):

I will just say that I like your notations:)